OpenStudy (anonymous):

Solving Quadratic Equations? 7(x-4)²-18=10?

jimthompson5910 (jim_thompson5910):

I'll get you started 7(x-4)^2 - 18 = 10 7(x-4)^2 = 10+18 7(x-4)^2 = 28 (x-4)^2 = 28/7 (x-4)^2 = 4